Circular No. 7003

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 C. Kouveliotou, Universities Space Research Association (USRA) at

Marshall Space Flight Center (MSFC), NASA; T. Strohmayer, Goddard Space Flight Center (GSFC), NASA; T. Takeshima, USRA at GSFC; J. H. Swank, GSFC; and P. Woods, University of Alabama in Huntsville, report on behalf of a larger collaboration: “We have observed SGR 1900+14 with the RXTE/PCA and RXTE/HEXTE, following the possible RXTE/ASM detection of this source (Remillard et al., IAUC 7002); reactivation of the source has been observed with BATSE/CGRO. The RXTE observation started on Aug. 28.55 UT and lasted for 5760 s. We detect a source with average persistent flux of 6 mCrab, which was strongly pulsed (roughly 23-percent peak-to-valley amplitude) at 5.16 s (not yet corrected to the barycenter). During the observation, multiple short-duration bursts were observed (< 0.125 s), one with intensity above 1.4 Crab. Analysis of scans carried out to localize the source is hampered by its variability. While no improvement on the ASM position has been obtained, the pulse period and the bursts identify the source as SGR 1900+14, the pulsing burst source observed in June 1998 (IAUC 7001), but with 10 times the persistent flux. No evidence was seen of the 89.17-s pulse period reported in data of a 1996 obervation of SGR 1900+14 (IAUC 6882). Additional observations are being carried out.”

C. Kouveliotou, USRA; G. J. Fishman, NASA/MSFC; and P. Woods and M. Kippen, UAH, communicate on behalf of the BATSE/CGRO team: “BATSE did not detect the burst of Aug. 27.432148 UT (Cline et al., IAUC 7002) due to earth occultation of the source. However, we have detected a dozen short (∼ 1-s) bursts starting on ∼ Aug. 27.4456 UT and continuing to the present. Several of these bursts have been intense (BATSE did not trigger because the trigger channels were set to 300–1000 keV). On Aug. 29.4282 UT, BATSE triggered (no. 7031) on a very intense event with fluence ∼ 7 × 10^{−6} erg cm^{−2} (25–300 keV); the event lasted roughly 3.5 s and exhibited two main peaks, after which weak evidence of pulses during its decay can be seen. We have reset the BATSE trigger energy channels to 25–100 keV in anticipation of further activity from SGR 1900+14.”

NOVA SAGITTARII 1998 Visual magnitude estimates by K. Hornoch, Lelekovice, Czech Rep.:

July 24.87 UT, 11.6; Aug. 9.87, 11.6; 10.88, 11.7; 25.83, 11.6; 30.81, 11.6.

1998 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green